OpenStudy (anonymous):

find and simplify the difference quotient: f(x+h)-f(x) / h (h is not zero for the given function)

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (anonymous):

dont even know where to start

OpenStudy (anonymous):

f(x+h)+3x^2?

OpenStudy (anonymous):

/ h

OpenStudy (anonymous):

What is \(f(x)\) defined as?

OpenStudy (anonymous):

oh sorry i left that out haha f(x) = -3x^2

OpenStudy (anonymous):

Alright. You have it right so far: \[\frac{f(x+h)-f(x)}{h}=\frac{f(x+h)+3x^2}{h}.\] Now, put \(x+h\) where \(x\) is in \(-3x^2.\)

OpenStudy (anonymous):

3(x+h)^2?

OpenStudy (anonymous):

\(-3(x+h)^2\), but yes. Now simplify the top of that fraction and see if anything else can be done.

OpenStudy (anonymous):

youre gonna have to write it out I cant picture it

OpenStudy (anonymous):

\[\frac{-3(x+h)^2+3x}{h}\] What can you do from here? Simplify \(-3(x+h)^2\). Then combine like terms. What do you get then?

OpenStudy (anonymous):

h^2+x^2+2xh?

OpenStudy (anonymous):

multiply that by -3?

OpenStudy (anonymous):

Yup. Then what?